A binary mixture of methanol-water containing 60 mole % methanol is charged to a batch still with two equilibrium stages operating at a pressure of 1 bar. The distillation is carried out until 60 % of the charge (on molar basis) has been taken as distillate product. A total condenser is used and the reflux is at boiling point. The still operates at a constant reflux ratio of 1.0. Stating clearly any assumption made, determine the mean composition of the overhead product.


1
Expert's answer
2021-11-25T01:27:07-0500

60+60= 120/2= 60%